OpenStudy (anonymous):

For the function P=20e^-0.18t.... What is the initial quantity? What is the growth rate in both decimal and percentage form? Is this a growth or decay? ***not sure on this one since it's a natural log right? :/ thanks :)

jimthompson5910 (jim_thompson5910):

What is the initial quantity?

OpenStudy (anonymous):

umm 20e? :/ not sure if the e is included?

jimthompson5910 (jim_thompson5910):

when t = 0, P = ???

OpenStudy (anonymous):

umm i'm not very sure... 20?

jimthompson5910 (jim_thompson5910):

e is just a number, it turns out that e = 2.71828 It's like pi in a way

jimthompson5910 (jim_thompson5910):

so if that 'e' bugs you, then replace it with 2.71828

OpenStudy (anonymous):

ohh okay.. so would the initial quantity be 54.3656.. so 54.37 ?

jimthompson5910 (jim_thompson5910):

plug in t = 0 and evaluate P = 20*e^(-0.18t)

jimthompson5910 (jim_thompson5910):

P = 20*e^(-0.18t) is the same as P = 20*2.71828^(-0.18t) since e = 2.71828 roughly

OpenStudy (anonymous):

ohh oops... i multiplied the wrong thing haha one sec... 20*0.835^7 ?

OpenStudy (anonymous):

sorry i meant 20*0.835^t

jimthompson5910 (jim_thompson5910):

P = 20*2.71828^(-0.18t) P = 20*2.71828^(-0.18*0) P = 20*2.71828^(0) P = ???

OpenStudy (anonymous):

P=20?

OpenStudy (anonymous):

because anything to the 0 power is 1 right? so 20x1=20?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

so 20 is the initial quantity?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

okay :) and i'm a bit confused on the growth rate part... does that include the t? and i'm a bit lost on how to find the actual rate :/

jimthompson5910 (jim_thompson5910):

Hint: solve for r (1+r)^t = e^(-0.18t) (1+r)^t = e^(-0.18) ^ t 1+r = e^(-0.18) r = ??

OpenStudy (anonymous):

-0.16 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

so that's the growth rate? -0.16 decimal form -16% percentage form? so that makes this decay?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

okay awesome! :) this one was confusing haha :P thank youuu!! :D

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

oh sorry one more thing... what's the difference between the continuous growth rate and the growth rate? :/ my question asks for continuous... but are they the same thing? :/

OpenStudy (anonymous):

@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

This is continuous decay since we're involving the 'e' term if we had something in the form P = a(1+r)^t then it would be either a normal growth problem or a normal decay problem

OpenStudy (anonymous):

ohh okay so -0.16 and 16% are still the continuous decay rate?

jimthompson5910 (jim_thompson5910):

yes that is correct

OpenStudy (anonymous):

oh okay awesome!! thank you!! :D

jimthompson5910 (jim_thompson5910):

sure thing